OpenStudy (anonymous):

Cos(2x)=1/2 how do u find the exact values

OpenStudy (amistre64):

the definition for cos is: adj/hyp of right triangle. so the adj side is 1; the hyp is 2; We can draw this, or you can just remember the angle that has a cos of 1/2...60 degrees; 2x = 60 x = 60/2 = 30

OpenStudy (amistre64):

there is a second angle with a cos=1/2... -60; or x = -30